I push a 10 kg cart up a frictionless ramp, which has an angle of 30 degrees to the horizontal. When I am done, the cart is 2 meters above the ground. It takes me 2 minutes to do this. What amount of energy did I use to do this task? How much power did it take?


1
Expert's answer
2021-02-22T16:04:16-0500
"W=mgh=(10)(9.8)(2)=196\ J\\P=\frac{W}{t}=\frac{196}{2\cdot60}=1.63\ W"